Find the centre of mass of rod(thin) whose linear density varies as k=a+bx where a and b are constants

Grade:11

1 Answers

Khimraj
6 years ago
xcm=(1/M)oL  x.dm
dm=λ(x)dx=(A+Bx).dx
M=0L λ(x).dx = oL(A+Bx).dx = AL+B(L2/2)
xcm=(1/M)oLx.(A+Bx).dx = (1/M)[(A/2)L2+(B/3)L3]
              [(A/2)L2+(B/3)L3]
Xcm =    ---------------------------------
                  AL+B(L2/2)   
 
                  AL+(2B/3)L2
Xcm =    ---------------------------------
                     2A+BL
 
Hope it clears. If you like answer then please approve it.